OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 49x^2 - 28x + 16 9a^2 - 30a + 25 25b^2 - 45b - 81 16x^2 - 24x - 9

OpenStudy (anonymous):

@SolomonZelman @mlb004

OpenStudy (campbell_st):

for a perfect square you have \[(a + b)^2 = a^2 + 2ab + b^2\] or \[(a-b)^2 = a^2 - 2ab + b^2\] so looking at the 2 general cases above you can eliminate the last 2 choices as they have the last terms as negatives... so for the 1st and 2nd which has a middle term that is double the product

OpenStudy (solomonzelman):

Which one do you think it is ?

OpenStudy (anonymous):

I would lean towards a

OpenStudy (solomonzelman):

I would not -:(

OpenStudy (solomonzelman):

B is more reliable.

OpenStudy (anonymous):

Thanks!